OpenStudy (anonymous):

find the exact value of tan22.5 using the half angle formula

OpenStudy (jdoe0001):

so... what's 22.5 * 2?

OpenStudy (anonymous):

45

OpenStudy (jdoe0001):

\(\bf tan(22.5^o)=tan\left(\frac{22.5^o\cdot 2}{2}\right)=tan\left(\frac{45^o}{2}\right)\) so there, use that in the half-angle identity for tangent :)

OpenStudy (anonymous):

okay at that point what is the next step to be taken

OpenStudy (jdoe0001):

using the trig half-angle identity :)

OpenStudy (jdoe0001):

there are about 3 for tangent, so pick either

OpenStudy (anonymous):

is it \[\sqrt{1-\cos45}/\sqrt{1+\cos45}\]

OpenStudy (jdoe0001):

yeap

OpenStudy (jdoe0001):

\(\bf tan(22.5^o)=tan\left(\frac{22.5^o\cdot 2}{2}\right)=tan\left(\frac{45^o}{2}\right)=\pm\sqrt{\cfrac{1-cos(45^o)}{1+cos(45^o)}} \)

OpenStudy (jdoe0001):

and that's a common known angle, and the cosine would be in your Unit Circle

OpenStudy (anonymous):

so cos 45 is \[\sqrt{2}/2\] how do i finish it im not great with the algebra

OpenStudy (jdoe0001):

ok... one sec

OpenStudy (anonymous):

k

OpenStudy (jdoe0001):

\(\bf tan\left(\frac{45^o}{2}\right)=\pm\sqrt{\cfrac{1-cos(45^o)}{1+cos(45^o)}}\\ \quad \\ \implies \pm\sqrt{\cfrac{1-\frac{\sqrt{2}}{2}}{1+\frac{\sqrt{2}}{2}}}\qquad \textit{let us do the inside first}\implies \cfrac{1-\frac{\sqrt{2}}{2}}{1+\frac{\sqrt{2}}{2}}\\ \quad \\ \cfrac{\frac{2-\sqrt{2}}{2}}{\frac{2+\sqrt{2}}{2}}\implies \frac{2-\sqrt{2}}{2}\cdot \frac{2}{2+\sqrt{2}}\implies \cfrac{2-\sqrt{2}}{2+\sqrt{2}}\\ \quad \\ \textit{multiplying by the conjugate of the denominator}\\ \quad \\ \cfrac{2-\sqrt{2}}{2+\sqrt{2}}\cdot \cfrac{2-\sqrt{2}}{2-\sqrt{2}}\implies \cfrac{(2-\sqrt{2})^2}{2^2-\sqrt{2^2}}\\ \quad \\ \cfrac{4-4\sqrt{2}+\sqrt{2^2}}{4-2}\implies \cfrac{4-4\sqrt{2}+2}{2}\implies \cfrac{\cancel{6}-\cancel{4}\sqrt{2}}{\cancel{2}}\\ \quad \\ 3-2\sqrt{2}\Leftarrow\textit{that's the inside, thus}\\ \quad \\ tan\left(\frac{45^o}{2}\right)\sqrt{3-2\sqrt{2}}\)

OpenStudy (jdoe0001):

hmm well \(\bf tan\left(\frac{45^o}{2}\right)=\pm\sqrt{3-2\sqrt{2}}\) that is

OpenStudy (anonymous):

thank you